Shift report 20010820 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Aug 20 2001 - 08:02:31 EDT

  • Next message: Flemming Videbaek: "new shift list"

    ------------------------------------------------------------
    Date: Mon Aug 20 08:02:31 2001, Shift: 00:00-08:00
    Supervisor: sjs, with: ah
    Log book#7 pages 138 - 139
    
    Shift Summary:
    ------------------------------------------------------------
    It was a dark and stormy evening when, 
    suddenly, the lights flickered....
    
    Lost:  Magnet control, tsir, sc, fb3.  HV tripped for
    BB, H1, H2, TOFW.
    
    RHIC went into a "severe weather shutdown" mode. The
    current estimate is for  several more hours before 
    operations resume.  No beam during shift.
    
    Found on-line documentation on how to reset sc.  Good work 
    whoever wrote this!  Reseting sc also cleared tsir error.
    
    Called MCR to reset magnet control computer.  Once reset,
    system recovered with all currents at 0 and status in
    Stby-fault state.  Still haven't figured out how to 
    get magnets running--looks like we might have to get help
    for this...
    
    fb3 error took somewhat longer to clear.  After failing to 
    find any documentation on how to handle this situation, AH
    scanned the previous shift reports for mention of fb3.  
    Fortunately, a reference was found that directed us to p 8 
    of Log 7, where the DAQ reboot procedure is discussed.  
    After several tries, this was found to work. A short 
    pulser test run was taken to test the daq using the fb3
    partition.
    
    Current situation:
    All systems look operational, but...
    Need to bring up fields in magnets.
    Need to set HV back to correct values.
    
    Should have plenty of time...RHIC operations still trying
    to recover systems that they lost...
    
    (and, from a mid-westerners perspective, it wasn't even that
    severe a storm...)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Aug 20 2001 - 08:03:26 EDT